# Vulnerability of LLMs to Vertically Aligned Text Manipulations

### v3

@@ 1 Introduction
Text classification is one of the most common tasks in Natural Language Processing (NLP), encompassing a wide range of applications, including sentiment analysis, harmful content detection, and spam filtering Minaee et al. (2021); Howard and Ruder (2018); Mirończuk and Protasiewicz (2018); Wei and Zou (2019); Yin et al. (2019). Since the introduction of the Transformer architecture Vaswani et al. (2017), models based on this architecture, such as BERT models Devlin et al. (2019); Sanh et al. (2020) and GPT models Radford et al. (2019); Brown et al. (2020); OpenAI et al. (2024), have achieved impressive performance in text classification Fields et al. (2024a); Rodrawangpai and Daungjaiboon (2022); Wang and Banko (2021). These models typically operate by either fine-tuning encoder-based Transformers Sun et al. (2020); Prottasha et al. (2022); González-Carvajal and Garrido-Merchán (2020); Qasim et al. (2022) or leveraging decoder-based models to generate text outputs directly Chae and Davidson (2023); Fields et al. (2024b); Milios et al. (2023). Their success across various text classification tasks has made them indispensable tools in NLP.
However, despite these advances, language models remain sensitive to certain types of input variation. Seemingly minor formatting changes, such as line breaks, punctuation marks, or word order, can significantly affect model outputs Sclar et al. (2024); Wang et al. (2023). A particularly intriguing example of this is vertical text formatting, where words are arranged vertically rather than horizontally. Although this format poses no challenge for human readers, it can confuse language models (shown in Figure 1). Meanwhile, vertical text finds practical applications in areas such as representing tree structures in computing, bypassing social media detection, and even educational games, such as Sudoku word and acrostic poetry. Therefore, understanding how language models interpret vertical text is essential for enhancing their utility.
As the capabilities of large language models (LLMs) such as the GPT series Radford et al. (2019); Brown et al. (2020); OpenAI et al. (2024) and Llama series Dubey et al. (2024); Touvron et al. (2023) continue to expand, evaluating their ability to understand vertical input becomes increasingly important. Although some of the research on input sensitivity has focused on encoder-based models such as BERT Devlin et al. (2019); Sanh et al. (2020), the behavior of LLMs when faced with unconventional input formats, such as vertical text, remains underexplored. Given that LLMs are increasingly being applied in critical areas such as content moderation, spam filtering, and misinformation detection, understanding their ability to handle atypical input is crucial to ensure robustness in real-world applications Rojas-Galeano (2024); Hu et al. (2024); Wang and Chang (2022).
@@ 2 Related Works
The introduction of the transformer architecture marked the beginning of the LLM era in natural language processing Vaswani et al. (2017). This evolution started with early models such as GPT-2 Radford et al. (2019) and GPT-3 Brown et al. (2020) and advanced to more semantically capable and safer models such as ChatGPT, GPT-4, and GPT-4o OpenAI et al. (2024). These models have been further refined through post-training techniques, including RLHF Ouyang et al. (2022) and DPO Rafailov et al. (2024). Throughout their development, concerns and research on the vulnerability and susceptibility of LLMs to various attacks have remained a key focus.
For example, in recent years, researchers have developed various methods to "jailbreak" LLMs, causing them to generate harmful content Dong et al. (2024); Perez and Ribeiro (2022); Chao et al. (2024); Deng et al. (2024). Such attacks clearly highlight the inherent vulnerabilities of LLMs. However, these vulnerabilities extend beyond this issue. Studies have shown that LLMs can produce significantly different outputs with similar input texts depending on content order (e.g., in multiple choice questions) Wang et al. (2023). Moreover, LLMs often show pronounced sensitivity to basic punctuation and line breaks, further illustrating their fragility Sclar et al. (2024).
Recent research reveals that vertically formatted text can severely impair the comprehension abilities of Transformer encoder-based models, such as BERT, causing significant performance degradation Rusert (2024); Devlin et al. (2019). This vulnerability, where straightforward text can substantially deceive language models, raises concerns about the potential impact of such formatting on contemporary LLMs. If LLMs are similarly susceptible to vertically formatted text, this could pose a threat to real-world applications.
@@ 3.1 Word Selection
In previous work, researchers used a greedy approach to assess the impact of each word on the prediction probability of the model, thus determining the importance of individual words in a sentence Rusert (2024). However, this method is time-consuming and resource-intensive for LLMs. In this paper, we address this issue by leveraging prompt-based LLMs as evaluators to identify and extract key words from the text.
@@ 3.2 Word Transformation
Model SST-2 CoLA ♠ QNLI Rotten Tomatoes Jigsaw Toxicity Original Vertical Original Vertical Original Vertical Original Vertical Original Vertical Closed-Source Models GPT-3.5 93.00 65.00 80.00 47.00 85.00 69.00 92.00 57.00 85.00 62.00 GPT-4 96.00 67.00 90.00 49.00 89.00 71.00 93.00 64.00 89.00 58.00 GPT-4o-mini 95.00 66.00 89.00 50.00 90.00 71.00 91.00 61.00 85.00 57.00 GPT-4o 95.00 68.00 87.00 47.00 90.00 70.00 90.00 65.00 91.00 60.00 Open-Source Models Llama3-8B 89.00 61.00 75.00 50.00 83.00 62.00 86.00 42.00 88.00 58.00 Llama3-70B 96.00 67.00 85.00 48.00 86.00 63.00 91.00 46.00 88.00 58.00 Llama3.1-8B 93.00 51.00 80.00 49.00 83.00 59.00 89.00 54.00 80.00 64.00 Llama3.1-70B 96.00 66.00 84.00 50.00 84.00 66.00 92.00 63.00 87.00 62.00 Gemma2-9B 88.00 60.00 83.00 51.00 78.00 60.00 86.00 58.00 85.00 53.00 Gemma2-27B 94.00 58.00 87.00 50.00 83.00 64.00 89.00 54.00 88.00 55.00 Qwen1.5-72B 95.00 63.00 81.00 52.00 85.00 66.00 93.00 57.00 88.00 63.00 Qwen2-72B 96.00 60.00 84.00 50.00 88.00 62.00 93.00 59.00 91.00 59.00
@@ 4.1 Datasets
SST-2 Socher et al. (2013): A benchmark dataset for sentiment classification, featuring labeled movie reviews.
CoLA Warstadt et al. (2019): A dataset for evaluating linguistic acceptability, containing English sentences labeled as grammatically acceptable or unacceptable.
QNLI Wang et al. (2019): Derived from the Stanford Question Answering Dataset (SQuAD), which evaluates whether a sentence answers a specific question.
Rotten Tomatoes Pang and Lee (2005): A dataset of movie reviews used for sentiment analysis, with each review classified as positive or negative.
Jigsaw Toxicity111https://www.kaggle.com/c/jigsaw-toxic-comment-classification-challenge cjadams et al. (2017): A dataset of online comments labeled for various types of toxic behavior.
@@ 4.3 Metric
where is an indicator function that is 1 if the condition is true and 0 otherwise, represents the true label, is the predicted label, and is the total number of samples.
@@ 4.4 Experiment Results
Specifically, for the CoLA dataset, the model’s classification accuracy drops by nearly 40 percentage points when two key words are input vertically. Given that the random prediction accuracy for a binary classification problem is 50%, this indicates that the model essentially loses its classification ability.
@@
For the other three datasets, the decrease in accuracy after implementing vertical input for different models mostly ranges from 25 to 40 percentage points, which is a significant drop and almost indicates that current LLMs nearly lose their natural language understanding ability when encountering vertical input, even when the manipulation involves only a few key words.
@@
| Pred. | + | 43 | 5 |
| - | 3 | 49 | |
Model SST-2 CoLA QNLI Rotten Tomatoes Jigsaw Toxicity Closed-Source Models GPT-3.5 w/ CoT 61.00 50.00 59.00 53.00 62.00 GPT-4 w/ CoT 66.00 51.00 68.00 60.00 56.00 GPT-4o-mini w/ CoT 68.00 50.00 76.00 64.00 56.00 GPT-4o w/ CoT 71.00 52.00 74.00 66.00 66.00 Open-Source Models Llama3-8B w/ CoT 59.00 49.00 64.00 53.00 55.00 Llama3-70B w/ CoT 66.00 46.00 59.00 55.00 62.00 Llama3.1-8B w/ CoT 53.00 51.00 62.00 56.00 63.00 Llama3.1-70B w/ CoT 68.00 47.00 65.00 62.00 65.00 Gemma2-9B w/ CoT 60.00 49.00 64.00 61.00 52.00 Gemma2-27B w/ CoT 61.00 51.00 64.00 57.00 53.00 Qwen1.5-72B w/ CoT 63.00 51.00 63.00 59.00 60.00 Qwen2-72B w/ CoT 62.00 52.00 64.00 60.00 61.00
Potential Threat of Mislabelling. According to Table 1, vertically formatted input text could reduce the accuracy of LLMs on sentiment and toxicity classification tasks by 30 to 40 percentage points. A more concerning issue emerges when examining the confusion matrix in Table 2 and Table 3: On the SST-2 dataset, inputting some words in a vertical format to the large language model reduces the classification accuracy for negative sentences from 91% to 24%. Similarly, on the Jigsaw Toxicity dataset, this approach decreases the model’s accuracy for classifying harmful text from 86% to 28%.
@@
Number of Vertical Words via Accuracy. In Figure 3, we choose 9 different LLMs to conduct a detailed experiment on the relationship between the number of vertically formatted words and the model’s classification accuracy.
@@ 5.1 Failure of CoT
In the above experiments, we have the LLMs directly output classification results for two types of text inputs. This method somewhat bypasses the model’s reasoning process, which may result in inaccuracies. To address this, we consider using a Chain of Thought (CoT) approach Wei et al. (2023) to determine if prompting the model to output reasoning steps can mitigate its vulnerability to vertically formatted text inputs. We incorporate "think step by step" into the prompt to encourage the LLMs to provide a step-by-step reasoning process, aiming to help the model better understand the intricacies of vertically formatted text inputs Kojima et al. (2023).
However, the experimental results in Table 4 show that even with the use of the CoT, the LLMs still fail to recognize that some words are input in a vertical format, despite generating a corresponding reasoning process. We can observe that after using CoT with LLMs on different classification tasks, the classification accuracy either increases or decreases compared to not using it, which differs significantly from our expectations. Even when there is an improvement, the increase in accuracy is minimal, generally less than or equal to 3 percentage points. On the Rotten Tomatoes dataset, although the two models from the Llama3 series show a noticeable improvement of around 10 percentage points after using CoT, their final classification accuracy remains extremely low, only about 55%. Unexpectedly, powerful models like GPT-3.5 and GPT-4 show reduced accuracy across four datasets despite using CoT.
@@ 6.1 Disordered Tokens
Due to the natural left-to-right and top-to-bottom writing order of text, the tokenizer used by LLMs encodes input text in this sequence. For instance, the tokenizer for the Llama3.1-8B model represents the word "vertical" as a single token when input horizontally. However, when input vertically, it converts into a token sequence of length 15 due to the spaces and line breaks. This causes the LLM to lose its understanding of the complete word.
@@ 6.2 Lost Attention
In Figure 7, we use text from sentiment classification as an example, the original text is “He appears miserable throughout as he swaggers through his scenes”. We observe that when words are input into the large language model in a horizontal format, the model assigns strong attention weights to the word "miserable", which signifies a negative sentiment, in relation to "negative". This indicates that the model recognizes "miserable" as conveying a negative emotion. However, when the word "miserable" is split and input in a vertical format, its components fail to establish a strong connection with "negative", ultimately leading to incorrect classification predictions by the model.
Similarly, in Figure 8, the original toxic text is "You are a talented idiot who never fails to surprise me", when we input the harmful word "idiot" normally, the large language model connects "toxic" with "idiot" and classifies the text as harmful. However, when we input "idiot" in a vertical format, the model loses its ability to establish a strong connection between "toxic" and the new tokens ("i", "d", "o" and "t") that make up "idiot", leading it to conclude that the input is not harmful.
@@ Appendix A Large Language Models
Gemma2: The next-generation open-source model from Google, released on June 27, 2024, as an improved version of Gemma, available in 2B, 9B, and 27B parameter configurations.
@@ Appendix B Implementation Details
For the GPT series models, we utilize the OpenAI API222https://openai.com/ for model invocation. During the word selection phase, we set top_p to 1.0 and temperature to 0.0 to ensure consistent word selection. In the text classification phase, we adjust top_p to 0.95 and keep temperature at 0.0 to maintain the reliability of the model’s output.
@@
For each dataset, we randomly select 100 test samples for experimentation. To ensure fairness, the number of samples for each label is evenly distributed in the selected test data, particularly for unbalanced datasets.
@@ Appendix D Does Explicit Prompt Work?
The findings indicate that although explicit prompts help LLMs mitigate the negative impact caused by vertical input to some extent, the improvement is quite small compared to the few-shot prompt. On the SST-2 dataset, GPT-3.5 improves its accuracy score by only 2 percentage points with an explicit prompt, whereas it gains 15 percentage points with a few-shot prompt. Similarly, for the GPT-4 model on the Rotten Tomatoes dataset, the improvement is only 8 percentage points with an explicit prompt, compared to 26 with a few-shot prompt.
@@
Input Format Dataset Accuracy(%) Original SST-2 96.67 QNLI 92.00 Rotten Tomatoes 97.33 Jigsaw Toxicity 95.00 Vertical SST-2 96.33 QNLI 92.00 Rotten Tomatoes 97.33 Jigsaw Toxicity 95.33
@@
To better understand the vulnerabilities of LLMs to vertical input compared to the human brain, we recruit three graduate-level students who are native English speakers to conduct experiments on four different datasets. Each student is presented with choices and required to select one, repeating this process twice—once before and once after the vertical input. Each time, the texts are shuffled to prevent memorization.
The experimental results of human classification are presented in Table 5, with accuracy computed as the average performance of the three students for each task. The results indicate that vertical input poses no challenge to human cognition, as evidenced by a minimal accuracy difference of just 0.33% across the two datasets — likely due to normal randomness rather than the input format. This demonstrates that while vertical input is easily comprehensible for humans, it remains a difficult problem for current LLMs, highlighting both the potential risks associated with relevant applications and the gap in format understanding between humans and LLMs.

### v4

@@ 1 Introduction
Text classification is one of the most common tasks in Natural Language Processing (NLP), encompassing a wide range of applications, including sentiment analysis, harmful content detection, and spam filtering (Minaee et al., 2021; Howard and Ruder, 2018; Mirończuk and Protasiewicz, 2018; Wei and Zou, 2019; Yin et al., 2019). Since the introduction of the Transformer architecture (Vaswani et al., 2017), models based on this architecture, such as BERT models (Devlin et al., 2019; Sanh et al., 2020) and GPT models (Radford et al., 2019; Brown et al., 2020; OpenAI et al., 2024), have achieved impressive performance in text classification (Fields et al., 2024a; Rodrawangpai and Daungjaiboon, 2022; Wang and Banko, 2021). These models typically operate by either fine-tuning encoder-based Transformers (Sun et al., 2020; Prottasha et al., 2022; González-Carvajal and Garrido-Merchán, 2020; Qasim et al., 2022) or leveraging decoder-based models to generate text outputs directly (Chae and Davidson, 2023; Fields et al., 2024b; Milios et al., 2023). Their success across various text classification tasks has made them indispensable tools in NLP.
However, despite these advances, language models remain sensitive to certain types of input variation. Seemingly minor formatting changes, such as line breaks, punctuation marks, or word order, can significantly affect model outputs (Sclar et al., 2024; Wang et al., 2023). A particularly intriguing example of this is vertical text formatting, where words are arranged vertically rather than horizontally. Although this format poses no challenge for human readers, it can confuse language models (shown in Figure 1). Meanwhile, vertical text finds practical applications in areas such as representing tree structures in computing, bypassing social media detection, and even educational games, such as Sudoku word and acrostic poetry. Therefore, understanding how language models interpret vertical text is essential for enhancing their utility.
As the capabilities of large language models (LLMs) such as the GPT series (Radford et al., 2019; Brown et al., 2020; OpenAI et al., 2024) and Llama series (Dubey et al., 2024; Touvron et al., 2023) continue to expand, evaluating their ability to understand vertical input becomes increasingly important. Although some of the research on input sensitivity has focused on encoder-based models such as BERT (Devlin et al., 2019; Sanh et al., 2020), the behavior of LLMs when faced with unconventional input formats, such as vertical text, remains underexplored. Given that LLMs are increasingly being applied in critical areas such as content moderation, spam filtering, and misinformation detection, understanding their ability to handle atypical input is crucial to ensure robustness in real-world applications (Rojas-Galeano, 2024; Hu et al., 2024; Wang and Chang, 2022).
@@ 2 Related Works
The introduction of the transformer architecture marked the beginning of the LLM era in natural language processing (Vaswani et al., 2017). This evolution started with early models such as GPT-2 (Radford et al., 2019) and GPT-3 (Brown et al., 2020) and advanced to more semantically capable and safer models such as ChatGPT, GPT-4, and GPT-4o (OpenAI et al., 2024). These models have been further refined through post-training techniques, including RLHF (Ouyang et al., 2022) and DPO (Rafailov et al., 2024). Throughout their development, concerns and research on the vulnerability and susceptibility of LLMs to various attacks have remained a key focus.
For example, in recent years, researchers have developed various methods to ‘jailbreak’ LLMs, causing them to generate harmful content (Dong et al., 2024; Perez and Ribeiro, 2022; Chao et al., 2024; Deng et al., 2024). Such attacks clearly highlight the inherent vulnerabilities of LLMs. However, these vulnerabilities extend beyond this issue. Studies have shown that LLMs can produce significantly different outputs with similar input texts depending on content order (e.g., in multiple choice questions) (Wang et al., 2023). Moreover, LLMs often show pronounced sensitivity to basic punctuation and line breaks, further illustrating their fragility (Sclar et al., 2024).
Recent research reveals that vertically formatted text can severely impair the comprehension abilities of Transformer encoder-based models, such as BERT, causing significant performance degradation (Rusert, 2024; Devlin et al., 2019). This vulnerability, where straightforward text can substantially deceive language models, raises concerns about the potential impact of such formatting on contemporary LLMs. If LLMs are similarly susceptible to vertically formatted text, this could pose a threat to real-world applications.
@@ 3.1 Word Selection
In previous work, researchers used a greedy approach to assess the impact of each word on the prediction probability of the model, thus determining the importance of individual words in a sentence (Rusert, 2024). However, this method is time-consuming and resource-intensive for LLMs. In this paper, we address this issue by leveraging prompt-based LLMs as evaluators to identify and extract key words from the text.
@@ 3.2 Word Transformation
Model SST-2 CoLA ♠ QNLI Rotten Tomatoes Jigsaw Toxicity Original Vertical Original Vertical Original Vertical Original Vertical Original Vertical Closed-Source Models GPT-3.5 GPT-4 GPT-4o-mini GPT-4o Open-Source Models Llama3-8B Llama3-70B Llama3.1-8B Llama3.1-70B Gemma2-9B Gemma2-27B Qwen1.5-72B Qwen2-72B
@@ 4.1 Datasets
SST-2 (Socher et al., 2013): A benchmark dataset for sentiment classification, featuring labeled movie reviews.
CoLA (Warstadt et al., 2019): A dataset for evaluating linguistic acceptability, containing English sentences labeled as grammatically acceptable or unacceptable.
QNLI (Wang et al., 2019): Derived from the Stanford Question Answering Dataset (SQuAD), which evaluates whether a sentence answers a specific question.
Rotten Tomatoes (Pang and Lee, 2005): A dataset of movie reviews used for sentiment analysis, with each review classified as positive or negative.
Jigsaw Toxicity111https://www.kaggle.com/c/jigsaw-toxic-comment-classification-challenge (cjadams et al., 2017): A dataset of online comments labeled for various types of toxic behavior.
@@ 4.3 Metric
where is an indicator function that is if the condition is true and otherwise, represents the true label, is the predicted label, and is the total number of samples.
@@ 4.4 Experiment Results
Specifically, for the CoLA dataset, the model’s classification accuracy drops by nearly percentage points when two key words are input vertically. Given that the random prediction accuracy for a binary classification problem is , this indicates that the model essentially loses its classification ability.
@@
For the other three datasets, the decrease in accuracy after implementing vertical input for different models mostly ranges from to percentage points, which is a significant drop and almost indicates that current LLMs nearly lose their natural language understanding ability when encountering vertical input, even when the manipulation involves only a few key words.
@@
| Pred. | + | ||
Model SST-2 CoLA QNLI Rotten Tomatoes Jigsaw Toxicity Closed-Source Models GPT-3.5 w/ CoT GPT-4 w/ CoT GPT-4o-mini w/ CoT GPT-4o w/ CoT Open-Source Models Llama3-8B w/ CoT Llama3-70B w/ CoT Llama3.1-8B w/ CoT Llama3.1-70B w/ CoT Gemma2-9B w/ CoT Gemma2-27B w/ CoT Qwen1.5-72B w/ CoT Qwen2-72B w/ CoT
Potential Threat of Mislabelling. According to Table 1, vertically formatted input text could reduce the accuracy of LLMs on sentiment and toxicity classification tasks by to percentage points. A more concerning issue emerges when examining the confusion matrix in Table 2 and Table 3: On the SST-2 dataset, inputting some words in a vertical format to the large language model reduces the classification accuracy for negative sentences from to . Similarly, on the Jigsaw Toxicity dataset, this approach decreases the model’s accuracy for classifying harmful text from to .
@@
Number of Vertical Words via Accuracy. In Figure 3, we choose different LLMs to conduct a detailed experiment on the relationship between the number of vertically formatted words and the model’s classification accuracy.
@@ 5.1 Failure of CoT
In the above experiments, we have the LLMs directly output classification results for two types of text inputs. This method somewhat bypasses the model’s reasoning process, which may result in inaccuracies. To address this, we consider using a Chain of Thought (CoT) approach (Wei et al., 2023) to determine if prompting the model to output reasoning steps can mitigate its vulnerability to vertically formatted text inputs. We incorporate ‘think step by step’ into the prompt to encourage the LLMs to provide a step-by-step reasoning process, aiming to help the model better understand the intricacies of vertically formatted text inputs (Kojima et al., 2023).
However, the experimental results in Table 4 show that even with the use of the CoT, the LLMs still fail to recognize that some words are input in a vertical format, despite generating a corresponding reasoning process. We can observe that after using CoT with LLMs on different classification tasks, the classification accuracy either increases or decreases compared to not using it, which differs significantly from our expectations. Even when there is an improvement, the increase in accuracy is minimal, generally less than or equal to percentage points. On the Rotten Tomatoes dataset, although the two models from the Llama3 series show a noticeable improvement of around percentage points after using CoT, their final classification accuracy remains extremely low, only about . Unexpectedly, powerful models like GPT-3.5 and GPT-4 show reduced accuracy across four datasets despite using CoT.
@@ 6.1 Disordered Tokens
Due to the natural left-to-right and top-to-bottom writing order of text, the tokenizer used by LLMs encodes input text in this sequence. For instance, the tokenizer for the Llama3.1-8B model represents the word ‘vertical’ as a single token when input horizontally. However, when input vertically, it converts into a token sequence of length 15 due to the spaces and line breaks. This causes the LLM to lose its understanding of the complete word.
@@ 6.2 Lost Attention
In Figure 7, we use text from sentiment classification as an example, the original text is “He appears miserable throughout as he swaggers through his scenes”. We observe that when words are input into the large language model in a horizontal format, the model assigns strong attention weights to the word ‘miserable’, which signifies a negative sentiment, in relation to ‘negative’. This indicates that the model recognizes ‘miserable’ as conveying a negative emotion. However, when the word ‘miserable’ is split and input in a vertical format, its components fail to establish a strong connection with ‘negative’, ultimately leading to incorrect classification predictions by the model.
Similarly, in Figure 8, the original toxic text is ‘You are a talented idiot who never fails to surprise me’, when we input the harmful word ‘idiot’ normally, the large language model connects ‘toxic’ with ‘idiot’ and classifies the text as harmful. However, when we input ‘idiot’ in a vertical format, the model loses its ability to establish a strong connection between ‘toxic’ and the new tokens (‘i’, ‘d’, ‘o’ and ‘t’) that make up ‘idiot’, leading it to conclude that the input is not harmful.
@@ Appendix A Large Language Models
Gemma2: The next-generation open-source model from Google, released on June 27, 2024, as an improved version of Gemma, available in B, B, and B parameter configurations.
@@ Appendix B Implementation Details
For the GPT series models, we utilize the OpenAI API222https://openai.com/ for model invocation. During the word selection phase, we set top_p to and temperature to to ensure consistent word selection. In the text classification phase, we adjust top_p to and keep temperature at to maintain the reliability of the model’s output.
@@
For each dataset, we randomly select test samples for experimentation. To ensure fairness, the number of samples for each label is evenly distributed in the selected test data, particularly for unbalanced datasets.
@@ Appendix D Does Explicit Prompt Work?
The findings indicate that although explicit prompts help LLMs mitigate the negative impact caused by vertical input to some extent, the improvement is quite small compared to the few-shot prompt. On the SST-2 dataset, GPT-3.5 improves its accuracy score by only percentage points with an explicit prompt, whereas it gains percentage points with a few-shot prompt. Similarly, for the GPT-4 model on the Rotten Tomatoes dataset, the improvement is only percentage points with an explicit prompt, compared to with a few-shot prompt.
@@
Input Format Dataset Accuracy(%) Original SST-2 QNLI Rotten Tomatoes Jigsaw Toxicity Vertical SST-2 QNLI Rotten Tomatoes Jigsaw Toxicity
@@
To better understand the vulnerabilities of LLMs to vertical input compared to the human brain, we recruit three graduate-level students who are native English speakers to conduct experiments on four different datasets. Each student is presented with choices and required to select one, repeating this process twice: once before and once after the vertical input. Each time, the texts are shuffled to prevent memorization.
The experimental results of human classification are presented in Table 5, with accuracy computed as the average performance of the three students for each task. The results indicate that vertical input poses no challenge to human cognition, as evidenced by a minimal accuracy difference of just across the two datasets, likely due to normal randomness rather than the input format. This demonstrates that while vertical input is easily comprehensible for humans, it remains a difficult problem for current LLMs, highlighting both the potential risks associated with relevant applications and the gap in format understanding between humans and LLMs.